OpenStudy (anonymous):

Write an equation in-slope intercept form for the line that passes through (-3,2) and is perpendicular to the line described by y=3/4x + 4

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

That I would use the reciprical -2/3 for the slope

OpenStudy (katrinakaif):

What aboutthe y-intercept of a perpendicular line?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

your reciprocal would not be -2/3 it would be -4/3

OpenStudy (anonymous):

yeah use the reciprocal of the slope

OpenStudy (anonymous):

-4/3

OpenStudy (katrinakaif):

Once you have determined your slope, use: (y-y1) = m (x-x1) as (x1,y1) is equal to (-3,2)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Yeah, typo.

OpenStudy (anonymous):

y=-4/3x-2?

OpenStudy (anonymous):

y - 2 = -4/3(x +3) y - 2 = -4x/3 - 4 y = -4x/3 - 2 yep :D

OpenStudy (anonymous):

So I need to write it in slope-intercept. What is that formula used?

OpenStudy (anonymous):

y = mx + b is slope intercept. m -> slope b -> y-intercept your equation is... y=-4/3x-2 -4/3 = m -> slope -2 = b -> y-intercept y = mx + b gives you the slope and y-intercept of a line, enabling easy graph. y-y1 = m(x-x1) is point-slope form, as it gives a point and the slope x+y + c = 0, where c = some number like 1,2,3,4, 14/6, etc. is the general equation for a line.